Solve for r if 2r+b=10 and 2r-3b=-6
Answer by jim_thompson5910(35256)

Subtract equation 2 from equation

2r+b=10
2r-3b=-6
----------
0r+4b = 16

4b = 16
b = 16/4
b = 4

Use b = 4 to find r.

2r+b=10
2r+4=10
2r = 10-4
2r = 6
r = 6/2
r = 3


So r = 3 is your answer